OpenStudy (anonymous):

The vapor pressure of benzene is 100mmHg at 26.1C and 400mmHg at 60.6C. What is the boiling point of benzene at 760mmHg?

OpenStudy (anonymous):

Sounds like a Clausius-Clapeyron problem to me.

OpenStudy (abb0t):

omg. This. Lol. Does this look familiar?: \[P = Ae^{\frac{ -∆H_{vap} }{ RT} }\] or \[\ln(P)= \frac{ -∆H_{vap} }{ RT } + C\]